# Slicing Wasserstein over Wasserstein via Functional Optimal Transport

### v1

@@ 5 Numerical Experiments
In this section, we aim to showcase the numerical properties and benefits of our sliced distances. We start with the 1d case. Drawing a connection between meta-measures and the so-called Gromov–Wasserstein (GW) distance, we consider a shape classification experiment from (Piening & Beinert, 2025a). Next, we compare our multidimensional sliced distance to the s-OTDD (Nguyen et al., 2025). Finally, we present applications from the evaluation of point cloud distributions and perceptual image analysis. For all these experiments, we employ trapezoidal integration weights $w_{r}$. We refer to Appendix C for further experiments and details.
@@ 5.1 Shape Classification via Local Distance Distributions
First, we repeat a shape classification experiment from (Piening & Beinert, 2025a) based on parametrizing shapes as so-called metric measure (mm-)spaces. A mm-space is a tuple $(\mathcal{X},d_{\mathcal{X}},\mu)$ consisting of a compact metric space $(\mathcal{X},d_{\mathcal{X}})$ and a measure $\mu\in\mathcal{P}_{2}(\mathcal{X})$. Modelling data as (finite) metric spaces allows for invariance to isometric transformations such as rotations, often desirable for shapes. Particularly, we may parametrize 2d shapes as point clouds with pairwise Euclidean distances and 3d shapes as triangular meshes with pairwise surface distances (Beier et al., 2022).
@@
Now, we can represent two mm-spaces as $\mu_{\mathbb{X}},\nu_{\mathbb{Y}}\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}))$ and compare them using the 1d WoW distance (Mémoli, 2011, ‘Third Lower Bound’ (TLB’)), another SW-based distance (Piening & Beinert, 2025a, ‘Sliced Third Lower Bound’ (STLB)) or a so-called energy distance (Sato et al., 2020, ‘Anchor Energy’ (AE)). Alternatively, we may use our SQW distance.
Repeating experiments from (Piening & Beinert, 2025a), we precompute pairwise distance matrices with respect to our distance, TLB, STLB, AE, and the GW distance. Then, we estimate the accuracy of a k-nearest neighbor (KNN) classification by assigning each test point to the majority class among its three nearest neighbors. We average the classification accuracy over 1000 random 25%/75% training/test splits on four datasets. Based on preprocessing from (Piening & Beinert, 2025a), we employ the ‘2D Shapes’ dataset (Beier et al. (2022), $N=50$), the ‘Animals’ dataset (Sumner & Popović (2004), $N=50$) and the ‘FAUST’ dataset (Bogo et al., 2014) with 500 (‘FAUST-500’) and 1000 (‘FAUST-1000’) vertices per shape. We set the kernel parameter $\sigma$ to 0.01, $R=10$ and $S=100$. We use the same integration grid and projection number for STLB. We display the results in Table 1, where we observe comparable performance across all distances and a runtime advantage of SQW and STLB for the large-scale FAUST-1000 dataset, in particular against the GW distance.
Distance 2D shapes Animals FAUST-500 FAUST-1000
Acc. (%) Time (ms) Acc. (%) Time (ms) Acc. (%) Time (ms) Acc. (%) Time (ms)
@@ 5.2 Optimal Transport Dataset Distance
Next, we consider a comparison with the OTDD (Alvarez-Melis & Fusi, 2020) and the s-OTDD (Nguyen et al., 2025). These metrics have been developed to quantify the similarities between labelled datasets in a model-agnostic manner. Such similarity metrics are especially important for applications in transfer learning. In this area, it has been shown empirically that the OTDD and s-OTDD display a strong correlation with the performance gap in transfer learning and classification accuracy in data augmentation, see (Alvarez-Melis & Fusi, 2020). To assess the suitability of our sliced metric as a drop-in replacement for the computationally costly OTDD, we repeat an experiment from (Nguyen et al., 2025). We randomly split MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009) to create subdataset pairs, each ranging in size from 500 to 1000, and compute the OTDD, the s-OTDD, and our DSW between subdataset pairs.
We plot the results of our distance and the s-OTDD against OTDD for 100 dataset splits in Figure  2, where we include the Pearson and the Spearman correlation coefficients between both sliced distances and the OTDD. As our metric is originally defined on $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ and OTDD and s-OTDD on $\mathcal{P}(\mathcal{Y}\times\mathcal{P}(\mathbb{R}^{d}))$, we compute OTDD and s-OTDD with the label metric on $\mathcal{Y}$ set to zero for comparability, effectively representing each dataset as an empirial measure in $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ and computing the OTDD via WoW. To be precise, each class-conditional distribution is modeled as $\mu_{i}\in\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d})$ and the distribution over class-conditional distributions becomes our meta-measure $\mu\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d}))$. Similar to (Nguyen et al., 2025), we estimate both DSW ($R=10$, $\sigma=0.1$) and the s-OTDD (with Radon features) with $S=10,000$ projections. Based on this setting, we employ the original default implementation for the s-OTDD and the ‘exact’ OTDD. Treating OTDD as our ground truth, we clearly see a stronger correlation between DSW and OTDD for all datasets.
@@
Figure 2: Scatter plots and correlations ($\uparrow$) between the s-OTDD and the OTDD (top) and our DSW and the OTDD (bottom) for MNIST (2(a)), FashionMNIST (2(b)), and CIFAR-10 (2(c)).
@@ 5.3 Comparing Distributions of Point Clouds
Looking at the number of target shapes $M$ in Subfigure 4(a) (Class: ‘bed’, $N=10$, $\sigma_{\text{Noise}}=0$, $n=m=50$), we see that all metrics successfully capture mode collapse, i.e., $M=1$, and decrease for a larger number of target shapes $M$. Unlike the plateauing WoW and DSW metrics, OT-NNA displays an undesired behavior by increasing for $M\geq N=10$, however. As for the random Gaussian perturbations in Subfigure 4(b) (Class: ‘sofa’, $N=M=10$, $n=m=50$), all three metrics increase with increasing noise. Whereas OT-NNA is more sensitive to small noise levels, the WoW and DSW metrics are more sensitive to high noise levels. Considering the point cloud resolution $m$ in Subfigure 4(c) (Class: ‘monitor’, $N=M=10$, $n=500$), OT-NNA seems inconsistent regarding the resolution. In contrast, the WoW and DSW metrics are higher for $m~\leq~100$ and plateau after a seemingly sufficient resolution has been reached. Overall, we see that WoW and DSW display similar behavior as the OT-NNA and offer the advantage of being unbounded metrics. Additionally, DSW ($S=10,000$, $R=50$, $\sigma=0.1$) offers computational advantages as it takes around 0.25 seconds for $M=N=10$ and $m=n=500$, where WoW takes about 4.5 seconds and OT-NNA about 8.5 seconds (on our CPU). This makes it especially suitable for high-resolution point clouds and large point cloud batches.
@@
Figure 4: OT-NNA, WoW, and our DSW between target and reference point cloud batches for varying numbers of shapes $M$ in target batch (4(a)), for Gaussian noise $\sigma_{\text{Noise}}$ for target points (4(b)) and varying target point cloud resolution $m$ (4(c)). Fixed reference values are marked in red.
@@ 5.4 Comparing Image Distributions via their Patch Distributions
Given the importance of OT for imaging, we conclude with an imaging experiment. Interestingly, OT is utilized on two levels in this area. On the one hand, it is used to compare two batches of images using pairwise Euclidean distances (Genevay et al., 2018). On the other hand, using OT as a distance between two individual images remains of relevance due to the disadvantages of Euclidean distances. Those methods represent images as 2D histograms (Beier et al., 2023; Geuter et al., 2025) or as patch distributions (Hertrich et al., 2022; Elnekave & Weiss, 2022; Flotho et al., 2025). Thus, a natural combination is the comparison of image batches using WoW (Dukler et al., 2019).
@@
To validate this approach, we consider synthetic $64\times 64$ texture images based on random Perlin noise (Perlin, 1985). This texture synthesis model is controlled by several parameters, among them the lacunarity and the persistence. Similar to our previous experiment, we initialize a reference meta-measure $\bm{\mu}$ over images represented as patch distributions according to some reference parameters and compare it to a target meta-measure $\bm{\nu}$ with varying lacunarity, respectively, persistence. For batch size $32$ and patch size $p=8$, both meta-measures are initialized according to $32$ random images, where each image is represented by $(64-7)\times(64-7)=3249$ uniformly weighted patches of dimension $8\times 8=64$. In Figure 6, we compare the average behavior over five runs of the standard Wasserstein distance with Euclidean cost between our reference and target images and our patch-based DSW distance ($\sigma=0.1$, $S=10,000$ projections, $R=10$), where the ‘ground truth’ lacunarity and persistence reference parameters are again marked in red.
We observe that both distances are minimized for the ‘true’ reference parameters. Still, our patch-based approach is more sensitive to parameter variation and better at discriminating between different batches, see also supplementary comparisons. Note that patch-based WoW calculation takes about 40 seconds, whereas our sliced distance merely requires about one second.
@@
Figure 6: Comparing synthetic texture image batches via Euclidean Wasserstein and our sliced patch-based distance based on varying ‘lacunarity’ (6(a)) and ‘persistence’ (6(b)). Both distances are minimized for ‘true’ parameters (red), but our sliced distance leads to clearer discrimination.
@@ 6 Conclusion
We introduce a general sliced OT framework for measures on arbitrary Banach spaces. Leveraging the isometry between 1d Wasserstein and $L_{2}([0,1])$, Gaussian-process–parametrized $L_{2}$-projections, and classical spherical slicing, we define the DSW distance between meta-measures, a well-posed, computationally efficient substitute for WoW. We prove that DSW minimization corresponds to WoW minimization for discretized meta-measures and demonstrate practical effectiveness on datasets, shapes, and images. On the practical side, future work could align DSW with the original OTDD by employing hybrid slicing (Nguyen & Ho, 2024) to extend DSW to $\mathcal{P}_{2}(\mathcal{Y}\times\mathcal{P}_{2}(\mathbb{R}^{d}))$ or integrate convolutional projections (Nguyen & Ho, 2022) similar to the s-OTDD. Also, one might employ our Banach slicing for infinite-dimensional generative models (Hagemann et al., 2025). On the theoretical side, it would be of interest to analyze further topological properties.
@@ Appendix A Non-Spherical Sliced Wasserstein Distance on Banach Spaces
Since every measure on $U$ has a unique characteristic function, see (Ledoux & Talagrand, 1991, § 2.1), we conclude $\mu=\nu$. ∎
@@
$c_{1}\operatorname{SW}(\mu,\nu\,;\,\xi)\leq\operatorname{SW}(\mu,\nu)\leq c_{2}\operatorname{SW}(\mu,\nu;\,\xi)\quad\forall\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}).$ (35)
@@
This follows directly from Proposition A.4 and the identity $\operatorname{SW}(\mu,\nu)=C_{d}\operatorname{SW}(\mu,\nu;\eta)$, see (Nadjahi et al., 2021, Prop. 1). ∎

### v2

@@ 5 Numerical Experiments
In this section, we aim to showcase the numerical properties and benefits of our sliced distances. We start with the 1d case. Drawing a connection between meta-measures and the so-called Gromov–Wasserstein (GW) distance, we consider a shape classification experiment from (Piening and Beinert, 2025a). Next, we compare our multidimensional sliced distance to the s-OTDD (Nguyen et al., 2025). Finally, we present applications from the evaluation of point cloud distributions and perceptual image analysis. For all these experiments, we employ trapezoidal integration weights $w_{r}$. We refer to Appendix C for further experiments and details [^2] Code: https://github.com/MoePien/slicing_wasserstein_over_wasserstein.
@@ 5.1 Shape Classification via Local Distance Distributions
First, we repeat a shape classification experiment from (Piening and Beinert, 2025a) based on parametrizing shapes as so-called metric measure (mm-)spaces. A mm-space is a tuple $(\mathcal{X},d_{\mathcal{X}},\mu)$ consisting of a compact metric space $(\mathcal{X},d_{\mathcal{X}})$ and a measure $\mu\in\mathcal{P}_{2}(\mathcal{X})$. Modelling data as (finite) metric spaces allows for invariance to isometric transformations such as rotations, often desirable for shapes. Particularly, we may parametrize 2d shapes as point clouds with pairwise Euclidean distances and 3d shapes as triangular meshes with pairwise surface distances (Beier et al., 2022).
@@
Now, we can represent two mm-spaces as $\mu_{\mathbb{X}},\nu_{\mathbb{Y}}\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}))$ and compare them using the 1d WoW distance (Mémoli, 2011, ‘Third Lower Bound’ (TLB’)), another SW-based distance (Piening and Beinert, 2025a, ‘Sliced Third Lower Bound’ (STLB)) or a so-called energy distance (Sato et al., 2020, ‘Anchor Energy’ (AE)). Alternatively, we may use our SQW distance.
Repeating experiments from (Piening and Beinert, 2025a), we precompute pairwise distance matrices with respect to our SQW distance, TLB, STLB, AE, and the GW distance. Then, we estimate the accuracy of a k-nearest neighbor (KNN) classification by assigning each test point to the majority class among its three nearest neighbors. We average the classification accuracy over 1000 random 25%/75% training/test splits on four datasets. Based on preprocessing from (Piening and Beinert, 2025a), we employ the ‘2D Shapes’ dataset (Beier et al. (2022), $N=50$), the ‘Animals’ dataset (Sumner and Popović (2004), $N=50$) and the ‘FAUST’ dataset (Bogo et al., 2014) with 500 (‘FAUST-500’) and 1000 (‘FAUST-1000’) vertices per shape. Additionally, we report results on a synthetic version of MNIST (LeCun et al., 1998) denoted ‘MNIST-2000’ ($N=100$, 5 classes, 2000 points), whose details are given in Appendix C.4.1. We set the kernel parameter $\sigma$ to 0.01, $R=10$ and $S=100$. We use the same integration grid and projection number for STLB. We display the results in Table 1, where we additionally report the mean runtime of a single distance computation. We observe comparable performance across all distances and a runtime advantage of SQW and STLB for the large-scale FAUST-1000 and MNIST-2000 dataset, in particular against the GW distance. See Appendix C.4.2 for further studies.
Distance 2D shapes Animals FAUST-500 FAUST-1000 MNIST-2000
Acc. (%) Time (ms) Acc. (%) Time (ms) Acc. (%) Time (ms) Acc. (%) Time (ms) Acc. (%) Time (ms)
@@
Based on MNIST-2000, we additionally perform a parameter study to analyze the impact of the projection number $S$, grid size $R$, and kernel parameter $\sigma$. We again classify our MNIST-2000 dataset via SQW, but with varying parameters. The accuracies are displayed in the table in Figure 1, where we observe only slight variations of the resulting accuracies. Additionally, we analyze the correlation between TLB and SQW in Figure 1. The SQW distance is meant to emulate the WoW distance on $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}))$. Since TLB is exactly the WoW distance between $\mu_{\mathbb{X}}$ and $\mu_{\mathbb{Y}}$, we would therefore hope for a linear relation between SQW and TLB. To analyze this relation, we produce a scatter plot displaying all SQW-TLB pairs on MNIST-2000 for three different parameter configurations. In addition, we compute the Pearson and Spearman correlation coefficients. Overall, we see a clear linear relation with correlations coefficients of $0.99$ or more. Changing the kernel parameter $\sigma$ does not seem to impact this relation, but increasing the number of projections $S$ reduces the variance and raises the correlation from $0.99$ to $1.0$.
$S$ $R$ $\sigma$ Acc (%)
$10^{2}$ $10^{1}$ $0.1^{2}$ 84.8
$10^{3}$ $10^{1}$ $0.1^{2}$ 84.7
$10^{3}$ $10^{2}$ $0.1^{2}$ 84.7
$10^{2}$ $10^{2}$ $0.1^{2}$ 84.8
$10^{2}$ $10^{1}$ $0.1^{3}$ 85.0
$10^{2}$ $^{1} 0.110$$^{1}$ 82.6
Refer to caption Refer to caption Refer to caption
Figure 1: Left: Impact of projection number $S$, grid size $R$, and kernel parameter $\sigma$ on MNIST-2000 SQW classification. Right: Scatter plots and correlation between TLB (1d WoW, horizontal axis) and SQW (vertical axis) for different parameters on MNIST-2000.
@@ 5.2 Optimal Transport Dataset Distance
Next, we consider a comparison with the OTDD (Alvarez-Melis and Fusi, 2020) and the s-OTDD (Nguyen et al., 2025). These metrics have been developed to quantify the similarities between labelled datasets in a model-agnostic manner. Such similarity metrics are especially important for applications in transfer learning. In this area, it has been shown empirically that the OTDD and s-OTDD display a strong correlation with the performance gap in transfer learning and classification accuracy in data augmentation, see (Alvarez-Melis and Fusi, 2020). To assess the suitability of our DSW distance as a drop-in replacement for the computationally costly OTDD, we repeat an experiment from (Nguyen et al., 2025). We randomly split MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009) to create subdataset pairs, each ranging in size from 500 to 1000, and compute the OTDD, the s-OTDD, and our DSW between subdataset pairs.
We plot the results of our DSW distance and the s-OTDD against OTDD for 100 dataset splits in Figure  3, where we include the Pearson and the Spearman correlation coefficients between both sliced distances and the OTDD. As our DSW distance is originally defined on $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ and OTDD and s-OTDD on $\mathcal{P}(\mathcal{Y}\times\mathcal{P}(\mathbb{R}^{d}))$, we compute OTDD and s-OTDD with the label metric on $\mathcal{Y}$ set to zero for comparability, effectively representing each dataset as an empirial measure in $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ and computing the OTDD via WoW. To be precise, each class-conditional distribution is modeled as $\mu_{i}\in\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d})$ and the distribution over class-conditional distributions becomes our meta-measure $\mu\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d}))$. Similar to (Nguyen et al., 2025), we estimate both DSW ($R=10$, $\sigma=0.1$) and the s-OTDD (with Radon features) with $S=10,000$ projections. Based on this setting, we employ the original default implementation for the s-OTDD and the ‘exact’ OTDD. Treating OTDD as our ground truth, we clearly see a stronger correlation between DSW and OTDD for all datasets. Additional experiments can be found in Appendix C.5.
@@
Figure 3: Scatter plots and correlations ($\uparrow$) between the s-OTDD and the OTDD (top) and our DSW (‘Ours’) and the OTDD (bottom) for MNIST (3(a)), FashionMNIST (3(b)), and CIFAR-10 (3(c)).
@@ 5.3 Comparing Distributions of Point Clouds
Looking at the number of target shapes $M$ in Subfigure 5(a) (Class: ‘bed’, $N=10$, $\sigma_{\text{Noise}}=0$, $n=m=50$), we see that all metrics successfully capture mode collapse, i.e., $M=1$, and decrease for a larger number of target shapes $M$. Unlike the plateauing WoW and DSW metrics, OT-NNA displays an undesired behavior by increasing for $M\geq N=10$, however. As for the random Gaussian perturbations in Subfigure 5(b) (Class: ‘sofa’, $N=M=10$, $n=m=50$), all three metrics increase with increasing noise. Whereas OT-NNA is more sensitive to small noise levels, the WoW and DSW metrics are more sensitive to high noise levels. Considering the point cloud resolution $m$ in Subfigure 5(c) (Class: ‘monitor’, $N=M=10$, $n=500$), OT-NNA seems inconsistent regarding the resolution. In contrast, the WoW and DSW metrics are higher for $m~\leq~100$ and plateau after a seemingly sufficient resolution has been reached. Overall, we see that WoW and DSW display similar behavior as the OT-NNA and offer the advantage of being unbounded metrics. Additionally, DSW ($S=10,000$, $R=50$, $\sigma=0.1$) offers computational advantages as it takes around 0.25 seconds for $M=N=10$ and $m=n=500$, where WoW takes about 4.5 seconds and OT-NNA about 8.5 seconds (on our CPU). This makes it especially suitable for high-resolution point clouds and large point cloud batches. Further studies are contained in Appendix C.6.
@@
Figure 5: OT-NNA, WoW, and our DSW (‘Ours’)between target and reference point cloud batches for varying numbers of shapes $M$ in target batch (5(a)), for Gaussian noise $\sigma_{\text{Noise}}$ for target points (5(b)) and varying target point cloud resolution $m$ (5(c)). Fixed reference values are marked in red.
@@ 5.4 Comparing Image Distributions via their Patch Distributions
Given the importance of OT for imaging, we conclude with an imaging experiment. Interestingly, OT is utilized on two levels in this area. On the one hand, it is used to compare two batches of images using pairwise Euclidean distances (Genevay et al., 2018). On the other hand, using OT as a distance between two individual images remains of relevance due to the disadvantages of Euclidean distances. Those methods represent images as 2D histograms (Beier et al., 2023; Geuter et al., 2025) or as patch distributions (Hertrich et al., 2022; Elnekave and Weiss, 2022; Flotho et al., 2025). Thus, a natural combination is the comparison of image batches using WoW (Dukler et al., 2019).
@@
To validate this approach, we consider synthetic $64\times 64$ texture images based on random Perlin noise (Perlin, 1985). This texture synthesis model is controlled by several parameters, among them the lacunarity and the persistence, see Appendix C.7. Similar to our previous experiment, we initialize a reference meta-measure $\bm{\mu}$ over images represented as patch distributions according to some reference parameters and compare it to a target meta-measure $\bm{\nu}$ with varying lacunarity, respectively, persistence. For batch size $32$ and patch size $p=8$, both meta-measures are initialized according to $32$ random images, where each image is represented by $(64-7)\times(64-7)=3249$ uniformly weighted patches of dimension $8\times 8=64$. In Figure 7, we compare the average behavior over five runs of the standard Wasserstein distance with Euclidean cost between our reference and target images and our patch-based DSW distance ($\sigma=0.1$, $S=10,000$ projections, $R=10$), where the ‘ground truth’ lacunarity and persistence reference parameters are again marked in red.
We observe that both distances are minimized for the ‘true’ reference parameters. Still, our patch-based approach is more sensitive to parameter variation and better at discriminating between different batches, see also supplementary comparisons. Note that patch-based WoW calculation takes about 40 seconds, whereas our DSW distance merely requires about one second.
@@
Figure 7: Comparing synthetic texture image batches via Euclidean Wasserstein and our sliced patch-based distance based on varying ‘lacunarity’ (7(a)) and ‘persistence’ (7(b)). Both distances are minimized for ‘true’ parameters (red), but our DSW (‘Ours’) distance leads to clearer discrimination.
@@ 6 Conclusion
We introduce a general sliced OT framework for measures on arbitrary Banach spaces. Leveraging the isometry between 1d Wasserstein and $L_{2}([0,1])$, Gaussian-process–parametrized $L_{2}$-projections, and classical spherical slicing, we define the DSW distance between meta-measures, a well-posed, computationally efficient substitute for WoW. We prove that DSW minimization corresponds to WoW minimization for discretized meta-measures and demonstrate practical effectiveness on datasets, shapes, and images.
On the practical side, future work could align DSW with the original OTDD by employing hybrid slicing (Nguyen and Ho, 2024) to extend DSW to $\mathcal{P}_{2}(\mathcal{Y}\times\mathcal{P}_{2}(\mathbb{R}^{d}))$ or integrate convolutional projections (Nguyen and Ho, 2022) similar to the s-OTDD. Also, one might employ our Banach slicing for infinite-dimensional generative models (Hagemann et al., 2025) or integrate dynamic transport into our framework to enable applications in generative flow matching similar to (Chapel et al., 2025).
On the theoretical side, it would be of interest to analyze further topological properties similar to Han (2023). Another interesting direction is the question of sample complexity: In high-dimensional Euclidean spaces, sliced Wasserstein distances require fewer samples to approximate a continuous probability measure than the Wasserstein distance (Nadjahi et al., 2020). As WoW distances require many samples to approximate a meta-measure (Catalano and Lavenant, 2024), it would be interesting to analyze if our DSW distance displays better sample complexity properties.
@@ Appendix A Non-Spherical Sliced Wasserstein Distance on Banach Spaces
Since every measure on $U$ has a unique characteristic function, see (Ledoux and Talagrand, 1991, § 2.1), we conclude $\mu=\nu$. ∎
@@
$c_{1}\operatorname{SW}(\mu,\nu\,;\,\xi)\leq{\operatorname{SW}(\mu,\nu;\eta)}=\operatorname{SW}(\mu,\nu)\leq c_{2}\operatorname{SW}(\mu,\nu;\,\xi)\quad\forall\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}).$ (35)
@@
This follows directly from Proposition A.4 and the identity $\operatorname{SW}(\mu,\nu)=\operatorname{SW}(\mu,\nu;\eta)$, see (Nadjahi et al., 2021, Prop. 1). ∎